OpenStudy (anonymous):

Find the points on the ellipse 100x^2+y^2=100 that are farthest away from the point (1,0)

OpenStudy (freckles):

You know the distance formula?

OpenStudy (freckles):

You want to maximize that using the points (1,0) and (x,y)

OpenStudy (anonymous):

ok i know y^2=(100-100x^2)

OpenStudy (anonymous):

so d^2=x^2 + (100-100x^2)^2

OpenStudy (anonymous):

but im not sure about the x^2 part is it (1-x)^2 because the point is (1,0)?

OpenStudy (anonymous):

or am I just doing this all wrong? :\

OpenStudy (freckles):

\[d=\sqrt{(\Delta x)^2+(\Delta y)^2}=\sqrt{((x-1)^2+(y-0)^2}=\sqrt{(x-1)^2+y^2}\] I used the points given (1,0) and (x,y)

OpenStudy (freckles):

Okay and you squared the distance! Good job! :)

OpenStudy (freckles):

\[d^2=(x-1)^2+y^2\] Replace y^2 with what you found :)

OpenStudy (anonymous):

yep i did that ! :o

OpenStudy (freckles):

\[100x^2+y^2=100 => y^2=100-100x^2\]

OpenStudy (freckles):

Why did you square y^2?

OpenStudy (freckles):

\[d^2=(x-1)^2+100-100x^2\]

OpenStudy (anonymous):

OOOHH omg hahaahah thats where i went wrong?!

OpenStudy (anonymous):

so then after i take the derivative is do i find the roots ? -199x-1=0? that doesnt seem right :s

OpenStudy (freckles):

Nope nope.... So when you found ((x-1)^2)' You used chain rule right?

OpenStudy (freckles):

We know that (100)'=0

OpenStudy (freckles):

(-100x^2)'=-100(x^2)' we know to use power rule here

OpenStudy (anonymous):

yes so i get 2(x-1)*1 + 0 -200x

OpenStudy (freckles):

You need to distribute that 2. 2(x-1)=2x-1(2)=2x-2

OpenStudy (anonymous):

okay!

OpenStudy (freckles):

Right!: )

OpenStudy (freckles):

\[\text{ Set } (d^2)'=0\]

OpenStudy (freckles):

\[2x-2-200x=0\]

OpenStudy (anonymous):

ok but dont i have to divide by the 2d from when i differentiate (d^2)'

OpenStudy (freckles):

No no.

OpenStudy (anonymous):

why not ?

OpenStudy (freckles):

Minimizing d^2 will give us the same results as minimizing d

OpenStudy (anonymous):

ok, so i get x=1/98 ?

OpenStudy (anonymous):

then would that be my point? x= 1/98 and subbing that into the original equation y= 9.99948?

OpenStudy (anonymous):

also i only get 1 answer, shouldnt there be another point also?

OpenStudy (freckles):

\[y=\pm \sqrt{100-100x^2} \text{ correct?}\]

OpenStudy (freckles):

Say yes. lol

OpenStudy (anonymous):

but my x is only 1 point? and my y's will be 2 points? sorry im just a little confused

OpenStudy (freckles):

\[(\frac{1}{99}, \sqrt{100-100(\frac{1}{99})^2}) \text{ and } (\frac{1}{99},-\sqrt{100-100(\frac{1}{99})^2})\]

OpenStudy (anonymous):

okay but why 1/99? i thought we found that 2x-2-200x=0 then wouldnt x be 2/-198 ?

OpenStudy (freckles):